OpenStudy (anonymous):

Help on how to solve this? (x^2 - 4x +4)/( x^2 + 10x + 25) x (x+5)/(x^2 +3x -10)

OpenStudy (pfenn1):

Is this the expression you are trying to solve?\[\frac{x^2 - 4x +4}{ x^2 + 10x + 25} \times \frac{x+5}{x^2 +3x -10}\]

OpenStudy (anonymous):

yes

OpenStudy (pfenn1):

Let's try to factor \[x^2-4x+4=(x+a)(x+b)\]Do you know how to do this?

OpenStudy (anonymous):

yea you foil right ? o.O

OpenStudy (pfenn1):

yeah but it is kind of a "reverse" foil.

OpenStudy (anonymous):

ohh okay

OpenStudy (anonymous):

(x-2)(x-2) ?

OpenStudy (pfenn1):

yep, so how about factoring \[x^2 +10x+25?\]

OpenStudy (anonymous):

(x+5)(x+5)

OpenStudy (pfenn1):

ok and the last one is \[x^2+3x-10\]

OpenStudy (anonymous):

(x+5)(x-2)

OpenStudy (anonymous):

would it be x-2/ x+5?

OpenStudy (pfenn1):

Great. Now let's put them all back into the original equation and get\[\frac{(x-2)^2}{(x+5)^2}\times \frac{x+5}{(x-2)(x+5)}\] Can you further simplify?

OpenStudy (pfenn1):

No, not x-2/x+5.

OpenStudy (anonymous):

x-2/ x+5 ^2?

OpenStudy (pfenn1):

yep\[\frac{x-2}{(x+5)^2}\]